OpenStudy (anonymous):

Solve for x: 2^(x+4) = 4^(x-2) A. 0 B. 1 C. 4 D. 6 E. 8

OpenStudy (anonymous):

wait r u suppose deal with jux the exponents 1st?

OpenStudy (anonymous):

\[2^{(x+4)} = 4^{(x-2) }\]\[2^{(x+4)} = (2^2)^{(x-2) }\]\[2^{(x+4)} = (2)^{2(x-2) }\]\[2^{(x+4)} = 2^{(2x-4) }\]

OpenStudy (anonymous):

When we are left with this we note that in order for both sides to be equal the exponents have to be equal

OpenStudy (anonymous):

go with @Romero

OpenStudy (anonymous):

you are then left with trying to solve tthis equation for x \[x+4=2x-4\]

OpenStudy (anonymous):

As hard tricky as this question is all you really have to do is solve that last equation I posted for x

OpenStudy (anonymous):

The reason we set them equal is because the exponents to be equal. That will make the statement true

OpenStudy (anonymous):

oh so all u need to do for such questions is to find a common sort of coefficient?

OpenStudy (anonymous):

so in this the answer is 8 right?

OpenStudy (anonymous):

how did you get 8?

OpenStudy (anonymous):

Actually yeah you are right. I got confused.

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

x+4=2x-4 4=x-4 x=8

OpenStudy (anonymous):

oh ok thanx